OpenStudy (anonymous):

use the communative property and associative properties of addition to write an expression equivalent to(y/8+90)+6x. the y/8 is a fraction, help please

OpenStudy (anonymous):

(y/8 +90) + 6x Commutative property = (90 +y/8) + 6x Associative property = y/8 + ( 90 +6x)